# Aligning Foundation Model Priors and Diffusion-Based Hand Interactions for Occlusion-Resistant Two-Hand Reconstruction

### v1

@@ Abstract
Two-hand reconstruction from monocular images faces persistent challenges due to complex and dynamic hand postures and occlusions, causing significant difficulty in achieving plausible interaction alignment. Existing approaches struggle with such alignment issues, often resulting in misalignment and penetration artifacts. To tackle this, we propose a novel framework that attempts to precisely align hand poses and interactions by synergistically integrating foundation model-driven 2D priors with diffusion-based interaction refinement for occlusion-resistant two-hand reconstruction. First, we introduce a Fusion Alignment Encoder that learns to align fused multimodal priors — keypoints, segmentation maps, and depth cues from foundation models — during training. This provides robust structured guidance, further enabling efficient inference without foundation models at test time while maintaining high reconstruction accuracy. Second, we employ a two-hand diffusion model explicitly trained to transform interpenetrated poses into plausible, non-penetrated interactions, leveraging gradient-guided denoising to correct artifacts and ensure realistic spatial relations. Extensive evaluations demonstrate that our method achieves state-of-the-art performance on InterHand2.6M, FreiHAND, and HIC datasets, significantly advancing occlusion handling and interaction robustness. Our code will be publicly released.
@@ 1 Introduction
Recent advancements in the related area have highlighted the benefits of foundation model priors and generative priors in improving 3D reconstruction tasks. For example, in the task of 3D human recovery, WHAM [20] utilizes 2D human keypoint model to extract motion features as priors. TRAM [23] further employs segmentation maps and depth maps to help to regress the kinematic body motion of a human. BUDDI [15] reconstructs two individuals in close proximity by utilizing the diffusion model as a prior. These show that vision foundation model-inferred structured 2D cues, e.g. keypoint, segmentation mask, and depth, provide valuable prior knowledge for 3D hand reconstruction. However, a challenging issue persists: the misalignment between 2D foundation model priors and 3D hand representations. Directly applying or finetuning 2D information features with foundation models would be non-trivial in transforming them seamlessly into accurate 3D hand structures, as 2D-3D feature alignment is inherently ambiguous. Beyond this misalignment, 3D reconstruction introduces an additional challenge: interaction consistency. In occlusion scenarios, one hand may block key regions of the other in 2D space, making it difficult to infer plausible inter-hand interactions. Existing methods, especially hand reconstruction approaches[12, 25, 8], would still suffer from spatial inconsistencies, unnatural interactions, and interpenetration artifacts due to potentially weakly constrained interaction priors.
Figure 1: Two-hand recovery on InterHand2.6M (1st, 3rd rows), Re:InterHand (4th, 5th rows), and In-the-Wild (2nd, 6th rows).
To overcome the above alignment issues, we propose a unified framework that integrates foundation model-driven 2D priors with diffusion-based interaction refinement. We progressively utilize these two to enhance the alignment performance from 2D to 3D spaces for occlusion-resistant two-hand reconstruction. Firstly, we leverage the exceptional performance of foundational models in 2D vision tasks and use them to generate 2D keypoints, segmentation maps, and depth maps as strong priors for providing well-structured 2D reference guidance. 2D keypoints provide accurate spatial information about hand joints in the image plane, segmentation maps offer a detailed prior of hand shapes and boundaries, and depth maps encode pixel-level depth cues that better capture the hands’ relative positioning in 3D space. We then introduce a Fusion Alignment Encoder that learns to align key points, segmentation, and depth priors from a suite of vision foundation models [5] during training. This encoder distills 2D prior information into the same latent feature space and aligns them together and allows us to waive the explicit use of the vision foundation model during inference. Subsequently, a hand regressor estimates hand parameters based on the fused features. Secondly, we employ a two-hand diffusion model to align and ensure interaction consistency by learning to correct interpenetration artifacts through gradient-guided denoising. Aligning with plausible interactions is particularly important in hand reconstruction since one hand could usually occlude important fingers of the other hand, leading to penetration issues in 3D space. The introduced 2D vision priors are still primarily in 2D space and would not be sufficient to deal with the occluded parts. Considering that diffusion models are exceptional at complementing missing 2D vision cues, we devise a two-hand diffusion model that explicitly incorporates the inter-hand penetration as a condition to enable direct learning of plausible interaction between hands. This would ensure spatially consistent and occlusion-aware hand reconstructions. Overall, our approach effectively aligns vision foundation model priors while simultaneously ensuring plausible hand interactions with our diffusion model, ensuring that two-hand reconstructions remain geometrically accurate and interaction-consistent, even under occlusions. As shown in Figure 1, we achieve robust hand reconstruction across various scenes and interactive poses.
@@
  • We propose the integration of supplementary 2D reference information as prior constraints and structured guidance in the 2D domain for 3D hand reconstruction. Concurrently, we introduce a fusion alignment encoder designed to learn the vision foundation model at the latent feature level, thereby enabling the exemption of the vision foundational model during the inference phase.
  • We train a two-hand diffusion model conditioned on the interpenetrating hands to directly model the interaction priors. We employ the two-hand generation model with penetration gradient guidance to address the penetration issue after two-hand estimation.
  • Our approach achieves state-of-the-art performance across multiple datasets. Extensive ablation experiments further validate the improvements in model performance brought by the alignment of 2D priors and the diffusion-model-enhanced hand interactions. Our qualitative experiments in real-world scenarios also demonstrate superior alignment and depth recovery capabilities.
@@ 2 Related work
With the introduction of some high-quality hand datasets, recovering 3D hand MANO [19] parameters from monocular input images has recently achieved remarkable advancements. For single-hand recovery, METRO [28] employs a convolutional neural network to extract a single global image feature and performs position encoding by repeatedly concatenating this image feature with the 3D coordinates of a mesh template. MeshGraphormer [9] introduces a graph-convolution enhanced transformer to effectively model both local and global interactions. AMVUR [3] proposes a probabilistic approach to estimate the prior probability distribution of hand joints and vertices. Zhou et al. [29] simplifies the process by decomposing the mesh decoder into a token generator and a mesh regressor, achieving high performance and real-time efficiency through a straightforward yet effective baseline. HaMeR [17] highlights the significant impact of scaling up to large-scale training data and utilizing high-capacity deep architectures for improving the accuracy and effectiveness of hand mesh recovery. For two-hand recovery, IntagHand [7] propose a GCN-based network to reconstruct two interacting hands from a single RGB image, featuring pyramid image feature attention (PIFA) and cross hand attention (CHA) modules to address occlusion and interaction challenges. InterWild [12] bridges MoCap and ITW samples for robust 3D interacting hands recovery in the wild by leveraging single-hand ITW data for 2D scale space alignment and using geometric features for appearance-invariant space. ACR [25] explicitly mitigates interdependencies between hands and between parts by leveraging center and part-based attention for feature extraction. 4DHands [8] handles both single-hand and two-hand inputs while leveraging relative hand positions using a transformer-based architecture with Relation-aware Two-Hand Tokenization (RAT) and a Spatio-temporal Interaction Reasoning (SIR) module. Although these methods have generally achieved competitive results in hand pose and shape reconstruction, their performance in finer details is still lacking.
2.2 Integrating additional information
@@
Recently, many studies have attempted to introduce additional reference information as guidance in visual tasks to achieve better performance. For example, in the task of human digitization, ECON [24] takes as input an RGB image and is conditioned on the rendered front and back body normal images. This strategy allows it to excel at inferring high-fidelity 3D humans in loose clothing and challenging poses. For text-to-image generation, ControlNet [27] has also successfully utilized different types of conditional inputs, such as sketches, depth maps, and segmentation maps. It has successfully achieved the generation of images aligned with these conditional guides using a pretrained text-to-image diffusion model. For the 3D human motion estimation task, WHAM [20] uses human 2D keypoints to extract motion features as inputs for both the Motion Decoder and Trajectory Decoder. This approach achieves more robust and stable 3D human motion estimates in global coordinates. Zuo et al. [31] captured interaction priors in the latent space of a VAE and applied them to interacting hand reconstruction, effectively estimating plausible hand poses. InterHandGen [6] trained a cascaded two-hand generation model, which serves as a generative prior to formulate a loss regularizer for addressing the challenge of two-hand reconstruction. In this paper, we attempt to acquire intermediate 2D hand information as 2D domain constraints and utilize a two-hand diffusion-based interaction prior to respectively address 2D and 3D alignment challenges in hand parameter estimation.

### v2

# Learning to Align and Refine: A Foundation-to-Diffusion Framework for Occlusion-Robust Two-Hand Reconstruction

@@ Abstract
Two-hand reconstruction from monocular images faces persistent challenges due to complex and dynamic hand postures and occlusions, causing significant difficulty in achieving plausible interaction alignment. Existing approaches struggle with such alignment issues, often resulting in misalignment and penetration artifacts. To tackle this, we propose a dual-stage Foundation-to-Diffusion framework that precisely align 2D prior guidance from vision foundation models and diffusion-based generative 3D interaction refinement to achieve occlusion-robust two-hand reconstruction. First, we introduce a lightweight fusion alignment encoder that aligns fused multimodal 2D priors like key points, segmentation maps, and depth cues from vision foundation models during training. This provides robust structured guidance, further enabling efficient inference without heavy foundation model encoders at test time while maintaining high reconstruction accuracy. Second, we implement a two-hand diffusion model explicitly trained to convert interpenetrated 3D poses into plausible, penetration-free counterparts. Through collision gradient-guided denoising, the model rectifies artifacts while preserving natural spatial relationships between hands. Extensive evaluations demonstrate that our method achieves state-of-the-art performance on InterHand2.6M, HIC, and FreiHAND datasets, significantly advancing occlusion handling and interaction robustness. Our code will be publicly released.
@@ 1 Introduction
Recent advancements in the related area have highlighted the benefits of foundation model priors and generative priors in improving 3D reconstruction tasks. For example, in the task of 3D human recovery, WHAM [20] utilizes 2D human key point model to extract motion features as priors. TRAM [23] further employs segmentation maps and depth maps to help to regress the kinematic body motion of a human. BUDDI [15] reconstructs two individuals in close proximity by utilizing the diffusion model as a prior. These show that vision foundation model-inferred structured 2D cues, e.g. key point, segmentation mask, and depth, provide valuable prior knowledge for 3D hand reconstruction.
However, directly applying foundation models to 3D two-hand reconstruction is challenging: fine-tuning large 2D encoders introduces substantial computational overhead and suffers from ambiguous 2D-to-3D feature alignment. Furthermore, under mutual hand occlusion conditions, 2D priors (such as keypoints, segmentation maps, and depth) can become unreliable or inconsistent. Although 3D generative priors (e.g., diffusion models) could offer interaction modeling, they require accurate initial alignment to observed data. Without being grounded in reliable 2D cues, they may generate implausible or unstable results. Consequently, we propose a Foundation-to-Diffusion paradigm that strategically leverages the complementary strengths of both technologies to improve 2D-3D alignment for two-hand reconstruction. Due to the lack of dedicated alignment optimization, existing two-hand reconstruction methods [12, 25, 8] often suffer from spatial inconsistencies, unnatural interactions, and interpenetration artifacts.
Figure 1: Two-hand recovery on InterHand2.6M (1st, 3rd columns), Re:InterHand (4th, 5th columns), and In-the-Wild (2nd, 6th columns).
To this end, we propose a unified framework that integrates foundation model-driven 2D priors followed by generative 3D interaction refinement. We progressively utilize these two to enhance the alignment performance from 2D to 3D spaces for occlusion-resistant two-hand reconstruction.
Firstly, we leverage the vision foundational models in 2D vision tasks and use them to generate 2D key points, segmentation maps, and depth maps as priors for providing well-structured 2D reference guidance. However, directly employing large foundation model encoders proves computationally prohibitive. To address this, we propose a lightweight fusion alignment encoder that learns to consolidate key point, segmentation, and depth priors from multiple vision foundation model encoders [5] during training. Our encoder effectively distills these heterogeneous 2D priors into a unified latent space while eliminating the need for explicit foundation model computations during inference.
Secondly, we employ a two-hand diffusion model to align and ensure interaction consistency by learning to correct interpenetration artifacts through gradient-guided denoising. The 2D vision priors remain fundamentally limited to 2D space, lacking the capability to handle occluded regions effectively. Considering that 3D generative models are exceptional at complementing missing 2D vision cues, we devise a two-hand diffusion model that explicitly incorporates the inter-hand penetration as a condition to enable direct learning of plausible interaction between hands. This would ensure spatially consistent and occlusion-aware hand reconstructions.
Overall, our approach effectively aligns vision foundation model priors while simultaneously ensuring plausible hand interactions with our diffusion model, ensuring that two-hand reconstructions remain geometrically accurate and interaction-consistent, even under occlusions. As shown in Figure 1, we achieve robust hand reconstruction across various scenes and interactive poses.
@@
  • We introduce a lightweight fusion alignment encoder to replace the heavy foundation model encoder for learning fused 2D prior features as prior constraints and structured guidance in 2D domain, eliminating the foundation model during inference.
  • To directly model interaction priors while addressing two-hand penetration issues in 3D space, we train a two-hand diffusion model conditioned on interpenetrating hands and employ it with penetration gradient guidance during generation.
  • Together, these form a novel Foundation-to-Diffusion paradigm for occlusion-aware, realistic two-hand reconstruction, achieving state-of-the-art results on InterHand2.6M, HIC, and FreiHAND with ablations confirming the benefits of 2D priors alignment and diffusion-enhanced interactions. Real-world experiments further demonstrate superior alignment and depth recovery.
@@ 2 Related Work
With the introduction of some high-quality hand datasets, recovering 3D hand MANO [19] parameters from monocular input images has recently achieved remarkable advancements.
@@
Single-Hand Recovery: METRO [28] employs a convolutional neural network to extract a single global image feature and performs position encoding by repeatedly concatenating this image feature with the 3D coordinates of a mesh template. MeshGraphormer [9] introduces a graph-convolution enhanced transformer to effectively model both local and global interactions. AMVUR [3] proposes a probabilistic approach to estimate the prior probability distribution of hand joints and vertices. Zhou et al. [29] simplifies the process by decomposing the mesh decoder into a token generator and a mesh regressor, achieving high performance and real-time efficiency through a straightforward yet effective baseline. HaMeR [17] highlights the significant impact of scaling up to large-scale training data and utilizing high-capacity deep architectures for improving the accuracy and effectiveness of hand mesh recovery. Two-Hand Recovery: IntagHand [7] propose a GCN-based network to reconstruct two interacting hands from a single RGB image, featuring pyramid image feature attention (PIFA) and cross hand attention (CHA) modules to address occlusion and interaction challenges. InterWild [12] bridges MoCap and ITW samples for robust 3D interacting hands recovery in the wild by leveraging single-hand ITW data for 2D scale space alignment and using geometric features for appearance-invariant space. ACR [25] explicitly mitigates interdependencies between hands and between parts by leveraging center and part-based attention for feature extraction. 4DHands [8] handles both single-hand and two-hand inputs while leveraging relative hand positions using a transformer-based architecture with Relation-aware Two-Hand Tokenization (RAT) and a Spatio-temporal Interaction Reasoning (SIR) module.
Although these methods have generally achieved competitive results in hand pose and shape reconstruction, their performance in finer details is still lacking.
2.2 Integrating Task-Related Prior
Recent advances have demonstrated that incorporating task-related prior knowledge can significantly enhance performance in various visual tasks.
@@ 2.2 Integrating Task-Related Prior
2D Priors: ECON [24] takes as input an RGB image and is conditioned on the rendered front and back body normal images in human digitization task. This strategy allows it to excel at inferring high-fidelity 3D humans in loose clothing and challenging poses. For text-to-image generation, ControlNet [27] has also successfully utilized different types of conditional inputs, such as sketches, depth maps, and segmentation maps. It has successfully achieved the generation of images aligned with these conditional guides using a pretrained text-to-image diffusion model. For the 3D human motion estimation task, WHAM [20] uses human 2D key points to extract motion features as inputs for both the Motion Decoder and Trajectory Decoder. This approach achieves more robust and stable 3D human motion estimates in global coordinates. Generative Prior: Zuo et al. [31] captured interaction priors in the latent space of a VAE and applied them to interacting hand reconstruction, effectively estimating plausible hand poses. InterHandGen [6] trained a cascaded two-hand generation model, which serves as a generative prior to formulate a loss regularizer for addressing the challenge of two-hand reconstruction.
In this paper, we attempt to acquire multimodal 2D hand priors as 2D domain constraints from foundation models and utilize a two-hand diffusion-based interaction prior to respectively address 2D and 3D alignment challenges in two-hand parameter estimation.